# Properties preserved by classes of Chu transforms

### v1

@@ 3.2. Universal-flow formulas and preservation under surjective Chu transforms
In this subsection we focus on surjective Chu transforms, which were considered in [31]. Discussing further ramifications of his Theorem 3.6 above, van Benthem makes an informal remark [31, §7.4]: “if we know that the $f$-map in a Chu transform is surjective, then we can add universal object quantifiers afor-$\forall a$ in the construction of flow formulas”. We formalize his remark by introducing the class of universal-flow formulas (Definition 3.7) and then proving the corresponding preservation result (Proposition 3.8).
@@
We follow the proof of Proposition 3.5, using the induction on the complexity of a given universal-flow formula $\varphi$. The new case is to suppose that $\langle X,r,A\rangle\models\forall\bm{v}^{\mathsf{p}}\varphi[\bm{v}^{\mathsf{p}},\bm{x},g(\bm{b})]$ for some $\bm{x}$ in $X$ and $\bm{b}$ in $B$. If $\bm{y}$ is in $Y$, by surjectivity there exists $\bm{x}^{\prime}\in X$ such that $\bm{y}=f(\bm{x}^{\prime})$. We then have that $\langle X,r,A\rangle\models\varphi[\bm{x}^{\prime},\bm{x},g(\bm{b})]$. This implies inductively that $\langle Y,s,B\rangle\models\varphi[\bm{y},f(\bm{x}),\bm{b}]$, from which we obtain $\langle Y,s,B\rangle\models\forall\bm{v}^{\mathsf{p}}\varphi[\bm{v}^{\mathsf{p}},f(\bm{x}),\bm{b}]$. ∎
@@
Otto’s theorem is in turn a generalization of Lyndon’s interpolation theorem [19] and uses the same notation of the positive and negative occurrence of a symbol in formula. For that definition, we do not allow connectives $\implies$ and $\iff$, which anyway do not figure in our definition of ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas. For a symbol $S$ of a language $L$ and a formula $\varphi$, we say that $S$ occurs positively within $\varphi$ if $S$ has an occurrence within $\varphi$ which is within the scope of an even number of negation symbols. The symbol $S$ occurs negatively within $\varphi$ if $S$ has an occurrence within $\varphi$ which is within the scope of an odd number of negation symbols. Note that a symbol $S$ can occur both positively and negatively in $\varphi$ [^4] We can see that the connectives $\implies$ and $\iff$ are excluded since they introduce hidden negation symbols that can change the parity of an occurrence.
@@
Theorem 3.9 (Otto [23, Theorem 1]).
@@
    Every relation symbol which occurs positively in $\chi$, occurs positively both in $\varphi$ and $\psi$.
@@
    Every relation symbol which occurs negatively in $\chi$, occurs negatively both in $\varphi$ and $\psi$.
@@
Applying the version of Theorem 3.9 without equality, Otto [23, Theorem 14] gives an alternative proof of van Benthem’s Theorem 3.6. From that application, we shall now deduce a preservation theorem for first-order universal-flow formulas.
@@
An $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-formula is preserved under surjective Chu transforms if and only if it is logically equivalent to a first-order universal-flow formula.
@@
The right-to-left implication follows already from Proposition 3.8. Conversely, suppose an $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-formula $\varphi$ is preserved under surjective Chu transforms. As a preliminary step, as noted in Remark 3.4, $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-atomic formulas of the type $\lnot(v^{\mathsf{p}}_{0}=v^{\mathsf{p}}_{1})$ are not preserved unless $f$ is injective, which is not required for surjective Chu transforms. Hence, up to logical equivalence, $\varphi$ does not contain any occurrences of such atomic formulas. Similarly for atomic formulas of the type $(v^{\mathsf{s}}_{0}=v^{\mathsf{s}}_{1})$, which are only preserved if $g$ is injective.
Next, following the proof of [23, Theorem 14], let $\mathbb{U}=\langle U,V\rangle$ consist of two unary predicates for the points sort and the states sort, respectively. Let $\varphi^{\mathbb{U}}$ be obtained from $\varphi$ by restricting each quantifier accordingly. Otto shows that, if $\varphi$ is preserved under restrictions in the second sort (which is certainly true for our $\varphi$), then $\varphi^{\mathbb{U}}$ is equivalent to a $\mathbb{U}$-relativized formula $\chi$ in which $V$ occurs only negatively.
The fact that $\chi$ is $\mathbb{U}$-relativized simply means that $\chi$ has a rewriting as an $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-formula: let us call this rewriting $\varphi^{\ast}$. Then, the fact that $V$ occurs only negatively in $\chi$ means that $\varphi^{\ast}$ does not admit existential quantification over the sort corresponding to $V$, which is exactly how we defined universal-flow formulas. ∎
@@ 3.3. Regressive-flow formulas and preservation under restrictions in the second sort
Gemignani [15] considered topological properties preserved by continuous functions. He did not define what a “property” is but, for our purpose, we shall take it to mean any ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-sentence.
@@
(1) (Gemignani [15]) A property $P$ is regressive iff for every topological space $\langle X,\tau\rangle$ and every coarser topology $\tau^{\prime}\subseteq\tau$ on $X$, if $\langle X,\tau\rangle$ satisfies $P$ then also $\langle X,\tau^{\prime}\rangle$ satisfies $P$.
(2) An ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-sentence $\varphi$ is regressive iff it is preserved under restrictions in the second sort.
@@
  • all atomic ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas are atomic regressive-flow formulas;
@@
Since restrictions in the second sort are in particular surjective, the inductive step is the same as the proof of Proposition 3.8. Additionally, we only need to check that inequalities of the type $\lnot(v_{0}^{\mathsf{p}}=v_{1}^{\mathsf{p}})$ and equalities of the type $(v_{0}^{\mathsf{s}}=v_{1}^{\mathsf{s}})$ are preserved, but this clearly holds by $f$ and $g$ being identity functions and by the logical convention of $=$ being interpreted as the true equality. ∎
@@
Gemignani [15, Proposition 1] claimed, among other things, that if a regressive property is in addition preserved by continuous bijective functions, then it is preserved by all continuous surjective functions. However, this is false because, for example, the property of not being $T_{0}$ is regressive and preserved by continuous bijective functions, but is not preserved by continuous surjective functions.
@@ 4. Preservation of compactness and inconsistency-flow formulas
It was shown in [14] that surjective Chu transforms preserve compactness of abstract logics. Furthermore, in [11] it was shown that dense Chu transforms preserve $\langle\kappa,\lambda\rangle$-compactness between abstract logics admitting complements and ${<}\kappa$-intersections. Motivated by these results, in this section we give a general preservation theorem for dense Chu transforms. We explain in detail how it applies to the preservation of compactness both in topology and in logic. In particular, we obtain the earlier results as corollaries.
@@
Let $\kappa\leq\lambda$ be infinite cardinals. A Chu space is $\langle\kappa,\lambda\rangle$-compact if it satisfies the following ($\mathcal{L}^{\mathrm{II}}_{{(\lambda^{<\kappa})}^{+},\lambda^{+}}$-sentence:
(eq-1) $\forall\langle v^{\mathsf{s}}_{\alpha}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha<\lambda\rangle\biggl{(}\exists v^{\mathsf{p}}\bigwedge_{\alpha<\lambda}\lnot R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\ \lor\bigvee_{Z\in{[\lambda]}^{<\kappa}}\forall v^{\mathsf{p}}\bigvee_{\alpha\in Z}R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\biggr{)}.$
@@
  • every open cover of $X$ of cardinality $\leq\lambda$ has a subcover of cardinality $<\kappa$.
Let $\langle M,\models,S\rangle$ be an abstract logic which is closed under negation; then the following are equivalent:
  • the Chu space $\langle M,\models,S\rangle$ is $\langle\kappa,\lambda\rangle$-compact,
  • given a set of sentences $T$ of cardinality $\leq\lambda$, if each subset of $T$ of cardinality $<\kappa$ has a model, then $T$ itself has a model.
@@
The formula in (eq-1) is a universal-flow formula. Hence, as a consequence of Proposition 3.8, we obtain the following.
@@
Concentrating on the sublogic relation, García-Matos and Väänänen proved the relevant instance of Corollary 4.3 in [14, Lemma 2.3]. Subsequently, Džamonja and Väänänen generalized that result to dense Chu transforms [11, Corollary 2.6]. This motivates us to characterize which formulas are preserved by dense Chu transforms and the relation of such formulas with preserving compactness. We naturally connect this with the notion of consistency.
@@
In the case of an abstract logic $\langle M,\models,S\rangle$, the formula $\operatorname{Con}[a]$ is indeed expressing the consistency of the sentence $a$, in the sense of $a$ having a model. In topological spaces, $\operatorname{Con}[a]$ simply means that the open set $a$ is non-empty. However, flow formulas cannot express inconsistency because formulas of the form $\lnot\exists{\bm{v}}^{\mathsf{p}}\varphi$ or $\forall{\bm{v}}^{\mathsf{p}}\psi$ are not in the class of flow formulas. So we shall define the class of “inconsistency-flow formulas” as the formulas which can express not only consistency, but also inconsistency. They are formally defined as follows.
@@
Example 4.6.
Compactness is not preserved under dense Chu transforms and, therefore, is not expressible in general as an inconsistency-flow formula. To see this, let $Y$ be an infinite set and $p\in Y$. Equip $Y$ with the particular point topology (see the book [29] by Lynn Steen and J. Arthur Seebach) defined as $\tau(Y)=\{X\subseteq Y\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}X=\emptyset\text{ or }p\in X\}$. Then, the inclusion function $\{p\}\to Y$ gives a dense Chu transform from a compact space to a non-compact space.
@@
Let $\kappa\leq\lambda$ be infinite cardinals. A Chu space is $\langle\kappa,\lambda\rangle$-absolutely closed if it satisfies the following ($\mathcal{L}^{\mathrm{II}}_{{(\lambda^{<\kappa})}^{+},\lambda^{+}}$-sentence:
(eq-2) $\forall\langle v^{\mathsf{s}}_{\alpha}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha<\lambda\rangle\Bigl{[}\exists v^{\mathsf{p}}\!\bigwedge_{\alpha<\lambda}\!\lnot R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\ \lor\!\!\bigvee_{Z\in{[\lambda]}^{<\kappa}}\!\!\forall v^{\mathsf{s}}\bigl{[}\lnot\operatorname{Con}(v^{\mathsf{s}})\lor\exists v^{\mathsf{p}}\bigl{(}R(v^{\mathsf{p}},v^{\mathsf{s}})\land\bigvee_{\alpha\in Z}\!R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\bigr{)}\bigr{]}\Bigr{]}.$
@@
  • every open cover of $X$ of cardinality $\leq\lambda$ has a subset of cardinality $<\kappa$ whose union is dense in $X$.
Let $\langle M,\models,S\rangle$ be an abstract logic which is closed under negation; then the following are equivalent:
  • the Chu space $\langle M,\models,S\rangle$ is $\langle\kappa,\lambda\rangle$-absolutely closed,
@@
Since the formula in (eq-2) is an inconsistency-flow formula, from Proposition 4.5 we obtain the following.
@@
  • If a Chu space is $\langle\kappa,\lambda\rangle$-absolutely closed and admits complements and ${<}\kappa$-intersections, then it is $\langle\kappa,\lambda\rangle$-compact.
@@
$\forall\langle v^{\mathsf{s}}_{\alpha}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha\in Z\rangle\Bigl{[}\forall v^{\mathsf{p}}\bigvee_{\alpha\in Z}R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\longrightarrow\forall v^{\mathsf{s}}\bigl{[}\lnot\operatorname{Con}(v^{\mathsf{s}})\lor\exists v^{\mathsf{p}}\bigl{(}R(v^{\mathsf{p}},v^{\mathsf{s}})\land\bigvee_{\alpha\in Z}R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\bigr{)}\bigr{]}\Bigr{]}.$
Furthermore, the following holds in any Chu space which admits complements and ${<}\kappa$-intersections:
$\forall\langle v^{\mathsf{s}}_{\alpha}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha\in Z\rangle\Bigl{[}\forall v^{\mathsf{p}}\bigvee_{\alpha\in Z}R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\longleftrightarrow\forall v^{\mathsf{s}}\bigl{[}\lnot\operatorname{Con}(v^{\mathsf{s}})\lor\exists v^{\mathsf{p}}\bigl{(}R(v^{\mathsf{p}},v^{\mathsf{s}})\land\bigvee_{\alpha\in Z}R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\bigr{)}\bigr{]}\Bigr{]}.$

### v3

@@ 3.2. Universal-flow formulas and preservation under surjective Chu transforms
In this subsection we focus on surjective Chu transforms, which were considered in [10]. Discussing further ramifications of his Theorem 3.6 above, van Benthem makes an informal remark [10, §7.4]: “if we know that the $f$-map in a Chu transform is surjective, then we can add universal object quantifiers $\forall a$ in the construction of flow formulas”. We formalize his remark by introducing the class of universal-flow formulas (Definition 3.7) and then proving the corresponding preservation result (Proposition 3.8).
@@
We follow the proof of Proposition 3.5, using the induction on the complexity of a given universal-flow formula $\varphi$. The new case is to suppose that $\langle X,r,A\rangle\models\forall\bm{v}^{\mathsf{p}}\varphi[\bm{v}^{\mathsf{p}},\bm{x},g(\bm{b})]$ for some $\bm{x}$ in $X$ and $\bm{b}$ in $B$. If $\bm{y}$ is in $Y$, by surjectivity there exists $\bm{x}^{\prime}\in X$ such that $f(\bm{x}^{\prime})=\bm{y}$. We then have that $\langle X,r,A\rangle\models\varphi[\bm{x}^{\prime},\bm{x},g(\bm{b})]$. This implies inductively that $\langle Y,s,B\rangle\models\varphi[\bm{y},f(\bm{x}),\bm{b}]$, from which we obtain $\langle Y,s,B\rangle\models\forall\bm{v}^{\mathsf{p}}\varphi[\bm{v}^{\mathsf{p}},f(\bm{x}),\bm{b}]$. ∎
@@
Otto’s theorem is in turn a strengthening of Lyndon’s interpolation theorem [21] and uses the same notation of positive and negative occurrences of a symbol in a formula. For that definition, we do not allow connectives $\rightarrow$ and $\leftrightarrow$, which anyway do not figure in our definition of $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas. For a symbol $S$ of a language $\mathcal{L}$ and a formula $\varphi$, we say that $S$ occurs positively in $\varphi$ if $S$ has an occurrence within $\varphi$ which is in the scope of an even number of negation symbols. The symbol $S$ occurs negatively in $\varphi$ if $S$ has an occurrence within $\varphi$ which is in the scope of an odd number of negation symbols. Note that a symbol can occur both positively and negatively in a formula. [^4] We can see that the connectives $\rightarrow$ and $\leftrightarrow$ are excluded since they introduce hidden negation symbols that can change the parity of an occurrence.
@@
Theorem 3.9 (Otto [24, Theorem 1]).
@@
    Every relation symbol of $\mathcal{L}$ which occurs positively in $\chi$, occurs positively in both $\varphi$ and $\psi$.
@@
    Every relation symbol of $\mathcal{L}$ which occurs negatively in $\chi$, occurs negatively in both $\varphi$ and $\psi$.
@@
Applying the version of Theorem 3.9 without equality, Otto [24, Theorem 14] gave an alternative proof of van Benthem’s Theorem 3.6. From that application, we shall now deduce a preservation theorem for first-order universal-flow formulas.
@@
Suppose $\varphi(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})$ is an $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-formula which is preserved under surjective Chu transforms. Then there exists a first-order universal-flow formula $\theta(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})$ such that $\varphi$ and $\theta$ are logically equivalent and, if $R$ occurs positively (negatively) in $\theta$, then $R$ occurs positively (negatively) in $\varphi$.
@@
As a preliminary step, as noted in Remark 3.4, $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-atomic formulas of the type $\lnot(v^{\mathsf{p}}_{0}=v^{\mathsf{p}}_{1})$ are not preserved unless $f$ is injective, which is not required for surjective Chu transforms. Hence, up to logical equivalence, $\varphi$ does not contain any occurrences of such atomic formulas. Similarly for atomic formulas of the type $(v^{\mathsf{s}}_{0}=v^{\mathsf{s}}_{1})$, which are only preserved if $g$ is injective.
Next, following the proof of [24, Theorem 14], let $\mathbb{U}=\langle U,V\rangle$ consist of two unary predicates for the points sort and the states sort, respectively. Let $\varphi^{\mathbb{U}}$ be obtained from $\varphi$ by relativizing each quantifier accordingly. Otto shows that, if $\varphi$ is preserved under restrictions in the second sort (which is certainly true for our $\varphi$), then $\varphi^{\mathbb{U}}$ is logically equivalent to a $\mathbb{U}$-relativized formula $\chi$ in which $V$ occurs only negatively, and each free variable of $\chi$ is free in $\varphi^{\mathbb{U}}$.
The fact that $\chi$ is $\mathbb{U}$-relativized implies that $\chi$ has a rewriting as an $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-formula: let us call this rewriting $\theta(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})$. Then, the fact that $V$ occurs only negatively in $\chi$ means that $\theta$ does not admit existential quantification over the sort corresponding to $V$, which is exactly how we defined universal-flow formulas. Finally, the additional syntactic information on positive and negative occurrences of $R$ in $\theta$ can be deduced from points (2) and (3) of Theorem 3.9, as in the proof of [24, Proposition 11]. ∎
@@ 3.3. Regressive-flow formulas and preservation under restrictions in the second sort
Gemignani [17] considered topological properties preserved by continuous functions. He did not define what a “property” is but, for our purpose, we shall take it to mean any $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-sentence.
@@
  1. (1)
    (Gemignani [17]) A property $P$ is regressive iff for every topological space $\langle X,\tau\rangle$ and every coarser topology $\tau^{\prime}\subseteq\tau$ on $X$, if $\langle X,\tau\rangle$ satisfies $P$ then also $\langle X,\tau^{\prime}\rangle$ satisfies $P$.
  2. (2)
    An $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-sentence $\varphi$ is regressive iff it is preserved under restrictions in the second sort.
@@
  • all atomic $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas are atomic regressive-flow formulas;
@@
Since restrictions in the second sort are in particular surjective, the inductive step is the same as the proof of Proposition 3.8. Additionally, we only need to check that inequalities of the type $\lnot(v_{0}^{\mathsf{p}}=v_{1}^{\mathsf{p}})$ and equalities of the type $(v_{0}^{\mathsf{s}}=v_{1}^{\mathsf{s}})$ are preserved, but this clearly holds by $f$ and $g$ being identity functions and by the logical convention of == being interpreted as the true equality. ∎
@@
Gemignani [17, Proposition 1] claimed, among other things, that if a regressive property is in addition preserved by continuous bijective functions, then it is preserved by all continuous surjective functions. However, this is false because, for example, the property of not being $T_{0}$ is regressive and preserved by continuous bijective functions, but is not preserved by continuous surjective functions.
@@ 4. Preservation of compactness and inconsistency-flow formulas
It was shown in [16] that surjective Chu transforms preserve compactness of abstract logics. Furthermore, in [13] it was shown that dense Chu transforms preserve $\langle\kappa,\lambda\rangle$-compactness between abstract logics admitting complements and $\kappa-$intersections. Motivated by these results, in this section we give a general preservation theorem for dense Chu transforms. We explain in detail how it applies to the preservation of compactness both in topology and in logic. In particular, we obtain the earlier results as corollaries.
@@
Let $\kappa\leq\lambda$ be infinite cardinals. A Chu space is $\langle\kappa,\lambda\rangle$-compact if it satisfies the following $\mathcal{L}^{\mathrm{II}}_{{(\lambda^{<\kappa})}^{+},\lambda^{+}}$-sentence:
(1) $\forall\langle v^{\mathsf{s}}_{\alpha}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha<\lambda\rangle\biggl(\exists v^{\mathsf{p}}\bigwedge_{\alpha<\lambda}\lnot R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\ \lor\bigvee_{Z\in{[\lambda]}^{<\kappa}}\forall v^{\mathsf{p}}\bigvee_{\alpha\in Z}R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\biggr).$
@@
  • every open cover of $X$ of cardinality $\leq\lambda$ has a subcover of cardinality $\kappa$.
Let $\langle M,\vDash,S\rangle$ be an abstract logic which is closed under negation; then the following are equivalent:
  • the Chu space $\langle M,\vDash,S\rangle$ is $\langle\kappa,\lambda\rangle$-compact,
  • given a set of sentences $T$ of cardinality $\leq\lambda$, if each subset of $T$ of cardinality $\kappa$ has a model, then $T$ itself has a model.
@@
The sentence in (1) is universal flow. Hence, as a consequence of Proposition 3.8, we obtain the following corollary.
@@
Concentrating on the sublogic relation, García-Matos and Väänänen proved the relevant instance of Corollary 4.3 in [16, Lemma 2.3]. Subsequently, Džamonja and Väänänen generalized that result to dense Chu transforms [13, Corollary 2.6]. This motivates us to characterize which formulas are preserved by dense Chu transforms and the relation of such formulas with preserving compactness. We naturally connect this with the notion of consistency.
@@
In the case of an abstract logic, the formula $\operatorname{Con}[a]$ expresses the satisfiability of the sentence $a$, in the sense of $a$ having a model. In topological spaces, $\operatorname{Con}[a]$ simply means that the open set $a$ is non-empty. However, flow formulas cannot express inconsistency because formulas of the form $\lnot\exists\bm{v}^{\mathsf{p}}\varphi$ or $\forall\bm{v}^{\mathsf{p}}\psi$ are not in the class of flow formulas. So we shall define the class of “inconsistency-flow” formulas as the formulas which can express not only consistency, but also inconsistency. They are formally defined as follows.
@@
Examples 4.6.
  1. (1)
    For every infinite cardinal $\kappa$, the topological property of having a dense subset of cardinality $\leq\kappa$ is expressible by the following $\mathcal{L}^{\mathrm{II}}_{\kappa^{+},\kappa^{+}}$-inconsistency-flow sentence:
    $\exists\langle v^{\mathsf{p}}_{\alpha}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha<\kappa\rangle\forall v^{\mathsf{s}}\biggl(\lnot\operatorname{Con}(v^{\mathsf{s}})\lor\bigvee_{\alpha<\kappa}R(v^{\mathsf{p}}_{\alpha},v^{\mathsf{s}})\biggr).$
  2. (2)
    The topological property of connectedness is expressible by the following $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-inconsistency-flow sentence:
    $\forall v^{\mathsf{s}}_{0}\forall v^{\mathsf{s}}_{1}\bigl(\lnot\operatorname{Con}(v^{\mathsf{s}}_{0})\lor\lnot\operatorname{Con}(v^{\mathsf{s}}_{1})\\ \lor\exists v^{\mathsf{p}}(R(v^{\mathsf{p}},v^{\mathsf{s}}_{0})\land R(v^{\mathsf{p}},v^{\mathsf{s}}_{1}))\lor\exists v^{\mathsf{p}}(\lnot R(v^{\mathsf{p}},v^{\mathsf{s}}_{0})\land\lnot R(v^{\mathsf{p}},v^{\mathsf{s}}_{1}))\bigr).$
  3. (3)
    Compactness is not preserved under dense Chu transforms and, therefore, is not expressible in general by an inconsistency-flow formula. To see this, let $Y$ be an infinite set and $p\in Y$. Equip $Y$ with the particular point topology (see the book [36] by Lynn Steen and J. Arthur Seebach) defined as $\tau(Y)=\{X\subseteq Y\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}X=\emptyset\text{ or }p\in X\}$. Then, the inclusion function $\{p\}\to Y$ gives a dense Chu transform from a compact space to a non-compact space.
@@
Let $\kappa\leq\lambda$ be infinite cardinals. A Chu space is $\langle\kappa,\lambda\rangle$-absolutely closed if it satisfies the following $\mathcal{L}^{\mathrm{II}}_{{(\lambda^{<\kappa})}^{+},\lambda^{+}}$-sentence:
(2) $\forall\langle v^{\mathsf{s}}_{\alpha}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha<\lambda\rangle\biggl[\exists v^{\mathsf{p}}\bigwedge_{\alpha<\lambda}\lnot R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\\ \lor\bigvee_{Z\in{[\lambda]}^{<\kappa}}\forall v^{\mathsf{s}}\bigl[\lnot\operatorname{Con}(v^{\mathsf{s}})\lor\exists v^{\mathsf{p}}\bigl(R(v^{\mathsf{p}},v^{\mathsf{s}})\land\bigvee_{\alpha\in Z}R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\bigr)\bigr]\biggr].$
@@
  • every open cover of $X$ of cardinality $\leq\lambda$ has a subset of cardinality $\kappa$ whose union is dense in $X$.
Let $\langle M,\vDash,S\rangle$ be an abstract logic which is closed under negation; then the following are equivalent:
  • the Chu space $\langle M,\vDash,S\rangle$ is $\langle\kappa,\lambda\rangle$-absolutely closed,
@@
Since the sentence in (2) is inconsistency flow, from Proposition 4.5 we obtain the following corollary.
@@
  • If a Chu space is $\langle\kappa,\lambda\rangle$-absolutely closed and admits complements and $\kappa-$intersections, then it is $\langle\kappa,\lambda\rangle$-compact.
@@
$\forall\langle v^{\mathsf{s}}_{\alpha}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha\in Z\rangle\biggl[\forall v^{\mathsf{p}}\bigvee_{\alpha\in Z}R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\longrightarrow\forall v^{\mathsf{s}}\bigl[\lnot\operatorname{Con}(v^{\mathsf{s}})\lor\exists v^{\mathsf{p}}\bigl(R(v^{\mathsf{p}},v^{\mathsf{s}})\land\bigvee_{\alpha\in Z}R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\bigr)\bigr]\biggr].$
Furthermore, the following holds in any Chu space which admits complements and $\kappa-$intersections:
$\forall\langle v^{\mathsf{s}}_{\alpha}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}\alpha\in Z\rangle\biggl[\forall v^{\mathsf{p}}\bigvee_{\alpha\in Z}R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\longleftrightarrow\forall v^{\mathsf{s}}\bigl[\lnot\operatorname{Con}(v^{\mathsf{s}})\lor\exists v^{\mathsf{p}}\bigl(R(v^{\mathsf{p}},v^{\mathsf{s}})\land\bigvee_{\alpha\in Z}R(v^{\mathsf{p}},v^{\mathsf{s}}_{\alpha})\bigr)\bigr]\biggr].$